OpenStudy (anonymous):

which sum or difference is equivalent to the following expression 5 - 3x over 5

OpenStudy (anonymous):

i dont know math but narry.. that is all ~*~*~

OpenStudy (anonymous):

1 + 3/5 1 + 3x/5 1 - 3x/5 25 - 15x

OpenStudy (anonymous):



OpenStudy (anonymous):

@whydoihavetosignup1 @wlarson @willie12345678 @Euler271 @emcrazy14 @EthanJumper @ellie67 @eblack99 @robtobey @RaphaelFilgueiras @rockstargirl17 @rawalter26

OpenStudy (anonymous):

can someone help?

OpenStudy (anonymous):

\[\frac{ 5 - 3x }{ 5 } = \frac{ 5 }{ 5 } - \frac{ 3x }{ 5 }\]

OpenStudy (anonymous):

so 1- 3x/5

OpenStudy (anonymous):

Yes.

Directrix (directrix):

I agree with the work and answer shown. @B.R